What mass of ethene (C2H4) must burn to produce 46.4 g of carbon dioxide?

The UNBALANCED equation is shown below:

C2H4 + O2 → CO2 + H2O
Answer #1

Answer:

The balanced equation for the combustion of C2H4 is

C2H4 + 3O2 -------> 2CO2 + 2H2O

Given mass of CO2=46.4 g and molar mass of CO2=44.01 gmol.

Moles of CO2=mass/molar mass=46.4 g/44.01 g/mol=1.0543 mol.

From the balanced equation, moles of C2H4 used = (1/2) x moles of CO2 produced=(1/2) x 1.0543 mol=0.527 mol.

Molar mass of C2H4=28.04 g/mol

Therefore mass of C2H4 required=moles x molar mass=0.527 mol x 28.05 g/mol=14.787 g.

Thanks and I hope you like it. Please let me know if you have any doubt.